OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of 1/2 . (1 point) y - 6 = 1/2 (x + 3) y = 1/2 x - 6 y + 3 = 1/2 (x - 6) x - 2y = 12

OpenStudy (anonymous):

@Chris911 @helpout3

OpenStudy (ankit042):

y -y1 = m(x-x1)

OpenStudy (anonymous):

@ankit042 how do u know?

OpenStudy (texaschic101):

point slope form : y - y1 = m(x - x1) slope(m) = 1/2 (6,-3) x1 = 6 and y1 = -3 now just sub that into the formula

OpenStudy (ankit042):

it is the standard form ....put the values and you will have your answer

OpenStudy (anonymous):

O okay so B!!!

OpenStudy (ankit042):

How is it B? you made a small mistake

OpenStudy (anonymous):

O crap :/

OpenStudy (ankit042):

lol try again and you should get your answer!

OpenStudy (anonymous):

i need help with my question

OpenStudy (anonymous):

Okay I think I'm pretty sure it's A

OpenStudy (anonymous):

for my question?

OpenStudy (ankit042):

Nooo are you shooting in the dark lol x1 = 6 y1= -3 and m= 1/2 now can you try

OpenStudy (anonymous):

I'm sorry I've never been good at this stuff, I'll try again :/

OpenStudy (ankit042):

np :) just concentrate and you will get it!

OpenStudy (anonymous):

@texaschic101 I just saw that you commented on this haha sorry and @ankit042 I did it again with the formula texaschic101 gave me and the answer is C!

OpenStudy (texaschic101):

correct..it is C......good job :)

OpenStudy (anonymous):

Thank you both so much :)